OpenStudy (timmyk14):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (timmyk14):

y = 4/3x + 8 y = 4/3x + 3 y = -3/4x + 8 y = -3/4x + 3

OpenStudy (anonymous):

Ok well first what do you know about two parallel lines?

OpenStudy (timmyk14):

@hilbertboy96